OpenStudy (anonymous):

Situation: A 15 gram sample of a substance that's a by-product of fireworks has a k-value of .1405. find the substance's half-life, in days. round to the nearest tenth

OpenStudy (anonymous):

I answered your other question

OpenStudy (anonymous):

\[f(t) = A*e^{-kt}\] if f(t)/A = .5: \[.5 = e^{-kt}; -kt = \ln (.5); t = \ln(.5)/-k = 4.9 days\]

OpenStudy (anonymous):

4.9 is the answer